OpenStudy (anonymous):

Twelve and seventeen-eighteenths minus 4 and two-thirds is equal to what number?

OpenStudy (anonymous):

a. 8 and five-sixths b. 8 and five-thirteenths c. 8 and five-eighteenths d. 3 and one-sixteenth

OpenStudy (yanasidlinskiy):

\[12\frac{ 17 }{ 18 } - 4\frac{ 2 }{ 3 }\] Convert it into an improper fraction. Like this: \[12\frac{ 17 }{ 18 } = \frac{ 12\times 18 +17 }{ 18} = \frac{ 233 }{ 18 }\]

OpenStudy (paki):

12(17/18) - 4(2/3) 233/18 - 14/3 233-84/18 149/18 8.27.... @adventuretimefreak7